# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v3

@@ 6 Conclusion
In this paper, we propose a novel approach called PALM, for continual test-time adaptation. Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e., by backpropagating the KL divergence between a uniform distribution and the softmax outputs. This effectively eliminates the need for pseudo-labels. For the parameters of these layers, selected based on the norm of the gradients, the learning rates are adjusted based on their sensitivity to domain shift. Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines. The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.
@@
Acknowledgments. We thank all the reviewers for their helpful comments and suggestions. This project was supported by a grant from the University of Texas at Dallas, Richardson, TX, USA.

### v4

@@ Conclusion
In this paper, we propose a novel approach called PALM, for continual test-time adaptation. Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e., by backpropagating the KL divergence between a uniform distribution and the softmax outputs. This effectively eliminates the need for pseudo-labels. For the parameters of these layers, selected based on the norm of the gradients, the LRs are adjusted based on their sensitivity to domain shift. Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines. The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.
Acknowledgments
@@
We thank all the reviewers for their helpful comments and suggestions. This project was supported by a grant from the University of Texas at Dallas, Richardson, TX, USA.